# Nonparametric bounds for vaccine effects in randomized trials

### v1

@@ 7 Numerical examples
We conducted a simulation study to examine the behavior of the nonparametric bounds under various settings. The simulation study objectives were: (1) to evaluate how the extent of broken blinding influences the width of the true bounds; (2) to assess the consequences of violations of the monotonicity assumptions, specifically (2a) Assumption 8(i) and (2b) Assumption 9(i), on the validity of the bounds; (3) to evaluate the consequences of applying bounds derived under an incorrect causal structure; and (4) to assess the performance of bounds estimated from a finite sample.
@@
The full numerical example, including code and reproducibility files, is available at https://github.com/xlrod1/vaccine_for_publication.
@@ 7.1 DGM
We generated a two-arm vaccine trial for a large population of size $n=10^{6}$ following the DGM described below. For each participant, we simulated $\{A,U\}$ and then the full set of potential outcomes $\{S^{a},B^{a,m},Y^{a,m}\}$ for $a,m\in\{0,1\}\times\{-1,0,1\}$. First, when $m\in\{0,1\}$, we set $B^{a,m}=m$ and simulated according to the model
@@
where $\mathrm{expit}(x)=\frac{e^{x}}{1+e^{x}}$. Here, $\mathbb{I}_{\{\cdot\}}$ denotes a binary switch used to control whether Assumptions 9(i) hold: with $\mathbb{I}_{u}=1$ and $\mathbb{I}_{u^{2}}=0$ the assumption is satisfied, while with $\mathbb{I}_{u}=0$ and $\mathbb{I}_{u^{2}}=1$ it is violated (note that $\mathbb{I}_{u}$ and $\mathbb{I}_{u^{2}}$ control $U$ variable type but also its association with $Y$). The coefficients $\gamma_{B^{1}}$ and $\gamma_{B^{0}}$ represent the effects of $B$ on $Y$ when $a=1$ and $a=0$, respectively. These coefficients are used to control whether Assumptions 8(i) hold: when both are positive the assumption is satisfied, whereas setting $\gamma_{B^{1}}>0$ and $\gamma_{B^{0}}<0$ corresponds to violation.
@@ 7.2 Simulation setup
Objective Figure Assumptions
LP-
Propositions
Monotonicity-
Propositions
(1) 3(d) 8(i) and 9(i) 6 7
(2a) 3(d) 8(i) 6 7
(2b) 3(d) 9(i) 6 7
(3) 3(c) and 3(d) 8(i) and 9(i) 4 5
(4a) 3(d) 8(i) and 9(i) 6 7
(4b) 3(a) 8(i) and 9(i) 4 5
@@ 7.3 Results
For Objective (1), Figure 4 displays the resulting bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ across different levels of the effect of AEs on $B$ ($\beta_{S}$). In all scenarios, the monotonicity-based bounds were narrower than those obtained using the LP-based method. As $\beta_{S}$ increased, the bounds for $\mathrm{VE}(0)$ and $\mathrm{VE}_{T}$ became wider, while the bounds for $\mathrm{VE}(1)$ became narrower. Importantly, even when $\beta_{S}=0$, the belief remained affected by $U$, which prevented point identification of the VEs.
Figure 4: The relationship between magnitude of broken blinding ($\beta_{S}$) and the width of the bounds for $VE(0)$, $VE(1)$ and $VE_{T}$ under the setting of Figure 3(d). Assumptions 8 and 9 are satisfied. A value of $\beta_{S}=0$ corresponds to the case where the belief was not influenced by AEs, $\beta_{S}=\log(1.5)$ represented moderate, and $\beta_{S}=\log(2)$ strong, broken blinding due to AEs. Circles represent the true values. LP: bounds constructed via Proposition 6 and the results in Appendix E; Monotonicity: bounds constructed via Proposition 7. Triangles at the lower boundary indicate bounds that extended below 0% but were truncated for display.
To address Objective (2a), Figure 5 presents the monotonicity-based bounds under violations of Assumption 9(i). For $\gamma_{U}>0$, as $\beta_{U}$ increased the monotonicity-based bounds became wider, whereas for $\gamma_{U}<0$ the pattern was mirrored – bounds became wider as $\beta_{U}$ decreased, with the effect most pronounced for $\mathrm{VE}(1)$ and $\mathrm{VE}_{T}$. Interestingly, when $\beta_{U}$ and $\gamma_{U}$ shared the same sign, the bounds remained valid despite the technical violation of monotonicity, while opposite signs led to clearly invalid bounds, particularly for $\mathrm{VE}_{T}$, as shown in Figure 5. This observation aligns with theoretical results from VanderWeele et al., (2008), who showed that the direction and magnitude of bias due to unmeasured confounding depend not only on whether monotonicity holds, but also on whether the confounder affects different pathways in opposing directions. When $\beta_{U}$ and $\gamma_{U}$ share the same sign, the effects of $U$ on $B$ and $Y$ are synergistic (both act in the same direction), preserving an approximately monotonic total effect of $U$. In contrast, when $\beta_{U}$ and $\gamma_{U}$ have opposite signs, the effects of $U$ on $B$ and $Y$ tend to cancel each other, creating non-monotonic behavior that violates Assumption 9(i).
@@
For Objectives (2b) and (3), Figures N.1N.3 examine the robustness of the bounds under different misspecifications. In some settings, particularly under stronger violations, the estimated lower bound exceeded the upper bound, indicating that the assumed causal structure or underlying assumptions (e.g., Assumption 8(i)) were incompatible with the DGM. In general, the monotonicity-based bounds were less robust to misspecification, as they were narrower. When Assumption 8(i) was violated, both LP-based and monotonicity-based bounds were robust for mild violation ($\gamma_{B1}=-\log(1.5)$) but became invalid for moderate to large violations ($\gamma_{B1}\in{-\log(3),-\log(5)}$) (Figure N.1). When the $S\rightarrow U$ path was incorrectly assumed to be absent, both the LP-based and monotonicity-based bounds remained valid for all values of $\delta_{U}$. However, the lower bound of the monotonicity-based interval for $\mathrm{VE}_{T}$ was very close to being invalid (Figure N.2). When both $S\rightarrow U$ and $S\rightarrow Y$ arrows were incorrectly assumed to be absent, only the LP-based bounds for $\mathrm{VE}(0)$ remained valid across all values of $\delta_{U}$, whereas all other bounds were either invalid or very close to becoming invalid (Figure N.3).
Finally, we turn to Objective (4). In practice, the bounds we developed in this paper must be estimated from finite samples. Tables N.1 and N.2 summarize the empirical coverage rates of the LP-based and monotonicity-based bounds. Overall, most intervals achieve coverage very close to the nominal 95% level, with only a few lower or upper limits showing slight under- or over-coverage, especially for the smaller sample size. For further discussion of the bootstrap CI coverage, see Section 9.
@@ 8 Real data example
The ENSEMBLE2 study is a randomized, double-blind, placebo-controlled, phase 3 trial of COVID-19 vaccines. We analyzed the data via the YODA platform (Krumholz and Waldstreicher,, 2016). The dataset available to us included a modestly larger cohort than the one reported in Hardt et al., (2022), since additional participants were enrolled after the primary analysis was completed. As a result, the numbers reported in this section slightly differ from those in Hardt et al., (2022). We considered semi-synthetic analysis to illustrate how this trial could have been analyzed and bounds could have been calculated had the belief was measured. The full code for the real data analysis is available at https://github.com/xlrod1/vaccine_for_publication.
@@
Since $B$ was not measured in this trial, we generated it to reflect two considerations. First, individuals who experienced AEs ($S=1$) were assumed to be more likely to believe they had received the vaccine. Second, we aimed to align $B$ with Assumption 8(i), which states that receiving a message of vaccination ($M=1$) increases the risk of infection ($Y=1$). As discussed in Section 4.2, this assumption implies that believing one is vaccinated ($B=1$) is also associated with a higher risk of infection ($Y=1$). Because $B$ was unobserved, we generated it in the opposite direction, individuals with a higher infection risk were made more likely to believe they were vaccinated. The full probabilities used to generate $B$ are provided in Appendix O.1.
@@
We first computed the point estimates of the VEs under Assumptions 13. We obtained $\widehat{VE(0)}=44.5\%$ (95% CI: 1.0%, 71.5%), $\widehat{VE(1)}=36.0\%$ (95% CI: -70.0%, 75.0%), and $\widehat{VE_{T}}=20.4\%$ (95% CI: -52.1%, 64.1%). Next, we estimated the LP-based and monotonicity-based bounds for $VE(0)$, $VE(1)$ and $VE_{T}$ plugging the sample proportions instead of the corresponding probabilities. Figure 6 presents the bounds under Figure 3(d), both with and without Assumption 8(i). The monotonicity-based bounds under Assumption 8(i) were the most informative. These bounds were notably narrower than the LP-based bounds and included the corresponding point estimates within their ranges. The bounds for $VE(0)$, (36.5%, 47.0%), suggest that the VE among participants who received the message $m=0$ could be even higher than the estimated $\widehat{VE(-1)}=39.3\%$ in the safety subset, reaching up to 47.0%. The bounds for $VE(1)$, (23.9%, 48.9%), indicate that the VE for those who received the message $m=1$ could be stronger than $VE(0)$ (and hence stronger than VE(1)VE(-1)), but could also be considerably lower than the point estimate $\widehat{VE(1)}=36.0\%$ and $\widehat{VE(-1)}$. Finally, the bounds for $VE_{T}$, (20.4%, 39.3%), suggest that the maximum total VE is similar to the published estimate, implying that the overall efficacy remains modest.
@@
Figure O.1 and Table O.2 present the LP-based and the monotonicity-based bounds under Figure 3(a), both with and without imposing Assumption 8(i). Under Assumption 8(i), in most cases we obtained upper bounds that were lower than the corresponding lower bounds, indicating that, under Figure 3(a), at least one of the Assumptions 8(i) or 9(i) must be violated. Under both Figures 3(a) and 3(d), when Assumption 8(i) was not imposed, both the LP-based and monotonicity-based bounds were wide and largely non-informative, with upper limits close to 100% and negative lower limits. This finding reflects that, without additional assumptions on how $M$ affects $Y$, we cannot determine whether the vaccine provides protection in this setting (Tables O.1O.2). All 95% bootstrap confidence intervals are reported in Table O.2 for Figure 3(a) and in Table O.1 for Figure 3(d).
@@
Figure 6: Estimated LP-based and monotonicity-based bounds under Figure 3(d). Each horizontal line shows the identified interval for $VE(0)$, $VE(1)$, and $VE_{T}$, stratified by method (LP vs. Monotonicity) and whether Assumption 8 (i) was imposed (solid = no, dashed = yes). Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display.
@@ 9 Discussion
In this paper, we address the challenge of disentangling the immunological and behavioral components of vaccine effects when the identification assumptions proposed by Stensrud et al., (2024) do not hold. Our work is motivated by growing evidence of compromised blinding in vaccine trials, such as the ENSEMBLE2 study, where unblinding is likely. In such cases, the standard VE estimate from the trial may not reflect the true immunological effect. However, the methodology proposed by Stensrud et al., (2024) to identify the immunological effect in such cases, is not applicable here, as the unobserved common cause may affect belief, side effects, and the outcome simultaneously, violating key assumptions.
We propose large-sample bounds using two complementary approaches: LP-based bounds and monotonicity-based bounds. We consider several plausible causal DAGs, focusing on four groups of scenarios: (i) when the side effect variable $S$ is unmeasured; (ii) when an unmeasured variable $U$ affects both belief and outcome; (iii) when $U$ affects belief, side effects, and outcome; and (iv) when two unmeasured variables influence the system. For each scenario, we derive appropriate bounds under the corresponding graphical model.
@@
To construct confidence intervals, we used percentile bootstrap-based intervals, as proposed in recent work applying LP bounds (Gabriel et al., 2023b, ; Gabriel et al., 2023a, ). However, under DAGs like Figure 3(a), where bounds involve the minimum and maximum of observed probabilities (observed up to estimation), standard bootstrap procedures may yield inconsistent intervals. As noted by Chernozhukov et al., (2013), such nonparametric bootstrap intervals can substantially underestimate the width of the true bounds, leading to undercoverage. This aligns with our simulations, where the empirical coverage of bootstrap intervals was often around 90% instead of 95%. While more sophisticated corrections and exact inference methods exist (Haile and Tamer,, 2003; Kreider and Pepper,, 2007; Chernozhukov et al.,, 2013), we did not apply these in the present study. As we have already pointed out, since the true DGM is unknown, we recommend computing the bounds under the structure shown in Figure 3(d), under which the bootstrap coverage was closest to 95%.
@@ Acknowledgments
The real data example section, carried out under YODA Project #2024-0972, used data obtained from the Yale University Open Data Access (YODA) Project, which has an agreement with Janssen Research & Development, L.L.C. The interpretation and reporting of research using this data are solely the responsibility of the authors and do not necessarily represent the official views of the Yale University Open Data Access Project or Janssen Research & Development, L.L.C. The original proposal is available at: https://yoda.yale.edu/data-request/2024-0972/.

### v2

@@ 7 Numerical examples
We conducted a simulation study to examine the behavior of the nonparametric bounds under various settings. The simulation study objectives were: (1) to evaluate how the extent of broken blinding influences the width of the true bounds; (2) to assess the consequences of violations of the monotonicity assumptions, specifically (2a) Assumption 101 and (2b) Assumption 81, on the validity of the bounds; (3) to evaluate the consequences of applying bounds derived under an incorrect causal structure; and (4) to assess the performance of bounds estimated from a finite sample.
@@ 7.1 DGM
We generated a two-arm vaccine trial following the DGM described below. To avoid sampling noise and approximate the true population values of the estimands and their bounds under the DGM, we simulated a large population of size $n=10^{6}$. For each participant, we simulated $\{A,U\}$ and then the full set of potential outcomes $\{S^{a},B^{a,m},Y^{a,m}\}$ for $a,m\in\{0,1\}\times\{-1,0,1\}$. First, when $m\in\{0,1\}$, we set $B^{a,m}=m$ and simulated according to the model
@@
where $\mathrm{expit}(x)=\frac{e^{x}}{1+e^{x}}$. Here, $\mathbb{I}_{\{\cdot\}}$ denotes a binary switch used to control whether Assumptions 10.1 hold: with $\mathbb{I}_{u}=1$ and $\mathbb{I}_{u^{2}}=0$ the assumption is satisfied, while with $\mathbb{I}_{u}=0$ and $\mathbb{I}_{u^{2}}=1$ it is violated (note that $\mathbb{I}_{u}$ and $\mathbb{I}_{u^{2}}$ control $U$ variable type but also its association with $Y$). The coefficients $\gamma_{B^{1}}$ and $\gamma_{B^{0}}$ represent the effects of $B$ on $Y$ when $a=1$ and $a=0$, respectively. These coefficients are used to control whether Assumptions 81 hold: when both are positive the assumption is satisfied, whereas setting $\gamma_{B^{1}}>0$ and $\gamma_{B^{0}}<0$ corresponds to violation.
@@ 7.2 Simulation setup
Objective Figure Assumptions LP- Propositions Monotonicity- Propositions
(1) 3(d) 8.1 and 10 5 6
(2a) 3(d) 8.1, 10.2 and 10.3 5 6
(2b) 3(d) 10 5 6
(3) 3(c) and 3(d) 81 and 10 3 4
(4a) 3(d) 8.1 and 10 5 6
(4b) 3(a) 8.1 and 10 3 4
@@ 7.3 Results
For Objective (1), Figure 4 displays the resulting bounds for $\mathrm{VE}(0)$, $\mathrm{VE}(1)$, and $\mathrm{VE}_{T}$ across different levels of the effect of $S$ on $B$ ($\beta_{S}$). In all scenarios, the monotonicity-based bounds were narrower than those obtained using the LP-based method. As $\beta_{S}$ increased, the bounds for $\mathrm{VE}(0)$ and $\mathrm{VE}_{T}$ became wider, while the bounds for $\mathrm{VE}(1)$ became narrower. Importantly, even when $\beta_{S}=0$, the belief remained affected by $U$, which prevented point identification of the VEs.
Figure 4: The relationship between magnitude of broken blinding ($\beta_{S}$) and the width of the bounds for $VE(0)$, $VE(1)$ and $VE_{T}$ under the setting of Figure 3(d). Assumptions 81 and 10 are satisfied. A value of $\beta_{S}=0$ corresponds to the case where the belief was not influenced by AEs, $\beta_{S}=\log(1.5)$ represented moderate, and $\beta_{S}=\log(2)$ strong, broken blinding due to AEs. Circles represent the true values. LP: bounds constructed via Proposition 5 and the results in Appendix E; Monotonicity: bounds constructed via Proposition 6. Triangles at the lower boundary indicate bounds that extended below 0% but were truncated for display.
To address Objective (2a), Figure 5 presents the monotonicity-based bounds under violations of Assumption 10.1. For $\gamma_{U}>0$, as $\beta_{U}$ increased the monotonicity-based bounds became wider, whereas for $\gamma_{U}<0$ the pattern was mirrored – bounds became wider as $\beta_{U}$ decreased, with the effect most pronounced for $\mathrm{VE}(1)$ and $\mathrm{VE}_{T}$. Interestingly, when $\beta_{U}$ and $\gamma_{U}$ shared the same sign, the bounds remained valid despite the technical violation of monotonicity, while opposite signs led to clearly invalid bounds, particularly for $\mathrm{VE}_{T}$, as shown in Figure 5. This observation aligns with theoretical results from VanderWeele et al. (2008), who showed that the direction and magnitude of bias due to unmeasured confounding depend not only on whether monotonicity holds, but also on whether the confounder affects different pathways in opposing directions. When $\beta_{U}$ and $\gamma_{U}$ share the same sign, the effects of $U$ on $B$ and $Y$ act in the same direction, preserving an approximately monotonic total effect of $U$. In contrast, when $\beta_{U}$ and $\gamma_{U}$ have opposite signs, the effects of $U$ on $B$ and $Y$ tend to cancel each other, creating non-monotonic behavior that violates Assumption 10.1.
@@
For Objectives (2b) and (3), Figures O.1O.3 examine the robustness of the bounds under different misspecifications. In some settings, particularly under stronger violations, the estimated lower bound exceeded the upper bound, indicating that the assumed causal structure or underlying assumptions (e.g., Assumption 8.1) were incompatible with the DGM. In general, the monotonicity-based bounds were less robust to misspecification, as they were narrower. When Assumption 8.1 was violated, both LP-based and monotonicity-based bounds were robust for mild violation ($\gamma_{B1}=-\log(1.5)$) but became invalid for moderate to large violations ($\gamma_{B1}\in{-\log(3),-\log(5)}$) (Figure O.1). When the $S\rightarrow U$ path was incorrectly assumed to be absent, both the LP-based and monotonicity-based bounds remained valid for all values of $\delta_{U}$. However, the lower bound of the monotonicity-based interval for $\mathrm{VE}_{T}$ was very close to being invalid (Figure O.2). When both $S\rightarrow U$ and $S\rightarrow Y$ arrows were incorrectly assumed to be absent, only the LP-based bounds for $\mathrm{VE}(0)$ remained valid across all values of $\delta_{U}$, whereas all other bounds were either invalid or very close to becoming invalid (Figure O.3).
Finally, we turn to Objective (4). In practice, the bounds we developed in this paper must be estimated from finite samples. Tables O.1 and O.2 summarize the empirical coverage rates of the LP-based and monotonicity-based bounds. Overall, most intervals achieve coverage very close to the nominal 95% level, with only a few lower or upper limits showing slight under- or over-coverage, especially for the smaller sample size. For further discussion of the bootstrap CI coverage, see Section 9.
@@ 8 Real data example
The ENSEMBLE2 study is a randomized, double-blind, placebo-controlled, phase 3 trial of COVID-19 vaccines. We analyzed the data via the YODA platform (Krumholz and Waldstreicher, 2016). The dataset available to us included a modestly larger cohort than the one reported in Hardt et al. (2022), since additional participants were enrolled after the primary analysis was completed. As a result, the numbers reported in this section slightly differ from those in Hardt et al. (2022). We considered semi-synthetic analysis to illustrate how this trial could have been analyzed and bounds could have been calculated had the belief was measured.
@@
Since $B$ was not measured in this trial, we generated it to reflect two considerations. First, individuals who experienced AEs ($S=1$) were assumed to be more likely to believe they had received the vaccine. Second, we aimed to align $B$ with Assumption 8.1, which states that receiving a message of vaccination ($M=1$) increases the risk of infection ($Y=1$). As discussed in Section 4.2, this assumption implies that believing one is vaccinated ($B=1$) is also associated with a higher risk of infection ($Y=1$). Because $B$ was unobserved, we generated it in the opposite direction, individuals with a higher infection risk were made more likely to believe they were vaccinated. The full probabilities used to generate $B$ are provided in Appendix P.1.
@@
We first computed the point estimates of the VEs under Assumptions 13. We obtained $\widehat{VE}(0)=44.5\%$ (95% CI: 1.0%, 71.5%), $\widehat{VE}(1)=36.0\%$ (95% CI: -70.0%, 75.0%), and $\widehat{VE_{T}}=20.4\%$ (95% CI: -52.1%, 64.1%). Next, we estimated the LP-based and monotonicity-based bounds for $VE(0)$, $VE(1)$ and $VE_{T}$ plugging the sample proportions instead of the corresponding probabilities. Figure 6 presents the bounds under Figure 3(d), both with and without Assumption 8.1. The monotonicity-based bounds under Assumption 8.1 were the most informative. These bounds were notably narrower than the LP-based bounds and included the corresponding point estimates within their ranges. The bounds for $VE(0)$, (36.5%, 47.0%), suggest that the VE among participants who received the message $m=0$ could be even higher than the estimated $\widehat{VE}(-1)=39.3\%$ in the safety subset, reaching up to 47.0%. The bounds for $VE(1)$, (23.9%, 48.9%), indicate that the VE for those who received the message $m=1$ could be stronger than $VE(0)$ (and hence stronger than VE(1)VE(-1)), but could also be considerably lower than the point estimate $\widehat{VE}(1)=36.0\%$ and $\widehat{VE}(-1)$. Finally, the bounds for $VE_{T}$, (20.4%, 39.3%), suggest that the maximum total VE is similar to the published estimate, implying that the overall efficacy remains modest.
@@
Figure P.1 and Table P.2 present the LP-based and the monotonicity-based bounds under Figure 3(a), both with and without imposing Assumption 8.1. Under Assumption 8.1, in most cases we obtained upper bounds that were lower than the corresponding lower bounds, indicating that, under Figure 3(a), at least one of the Assumptions 8.1 or 9.1 must be violated. Under both Figures 3(a) and 3(d), when Assumption 8.1 was not imposed, both the LP-based and monotonicity-based bounds were wide and largely non-informative, with upper limits close to 100% and negative lower limits. This finding reflects that, without additional assumptions on how $M$ affects $Y$, we cannot determine whether the vaccine provides protection in this setting (Tables P.1P.2). All 95% bootstrap confidence intervals are reported in Table P.2 for Figure 3(a) and in Table P.1 for Figure 3(d).
@@
Figure 6: Estimated LP-based and monotonicity-based bounds under Figure 3(d). Each horizontal line shows the identified interval for $VE(0)$, $VE(1)$, and $VE_{T}$, stratified by method (LP vs. Monotonicity) and whether Assumption 8 1 was imposed (solid = no, dashed = yes). Triangles at the left boundary indicate bounds that extended below 0% but were truncated for display.
@@ 9 Discussion
In this paper, we address the challenge of disentangling the immunological and behavioral components of vaccine effects when the identification assumptions proposed by Stensrud et al. (2024) do not hold. Our work is motivated by growing evidence of compromised blinding in vaccine trials, such as the ENSEMBLE2 study, where unblinding may be possible due to differences in AE rates. In such cases, the standard VE estimate from the trial may not reflect the true immunological effect. However, the methodology proposed by Stensrud et al. (2024) to identify the immunological effect in such cases is not applicable here, as the unobserved common cause may affect belief, AEs, and the outcome simultaneously, violating key assumptions.
We propose large-sample bounds using two complementary approaches: LP-based bounds and monotonicity-based bounds. We study three groups of plausible causal DAGs: (i) when $S$ is unmeasured; (ii) when $U$ affects both $B$ and $Y$; and (iii) when $U$ affects $B$, $S$, and $Y$, deriving bounds for each setting.
@@
To construct confidence intervals, we used percentile bootstrap-based intervals, as proposed in recent work applying LP bounds (Gabriel et al., 2023b, a). However, under DAGs like Figure 3(a), where bounds involve the minimum and maximum of observed probabilities (observed up to estimation), standard bootstrap procedures may yield inconsistent intervals. As noted by Chernozhukov et al. (2013), such nonparametric bootstrap intervals can substantially underestimate the width of the true bounds, leading to undercoverage. This aligns with our simulations, where the empirical coverage of bootstrap intervals was often around 90% instead of 95%. While more sophisticated corrections and exact inference methods exist (Haile and Tamer, 2003; Kreider and Pepper, 2007; Chernozhukov et al., 2013), we did not apply these in the present study. Since the true DGM is unknown, we recommend computing the bounds under the structure shown in Figure 3(d), under which the bootstrap coverage was closest to 95%.
@@
10 Software
Code and reproducibility materials for the numerical examples and the real-data analysis, with a synthetic dataset provided instead of the original data, are available at https://github.com/xlrod1/vaccine_for_publication.
@@ Acknowledgments
Section 8, carried out under YODA Project #2024-0972, used data obtained from the Yale University Open Data Access (YODA) Project, which has an agreement with Janssen Research & Development, L.L.C. The interpretation and reporting of research using this data are solely the responsibility of the authors and do not necessarily represent the official views of the Yale University Open Data Access Project or Janssen Research & Development, L.L.C. The original proposal is available at: https://yoda.yale.edu/data-request/2024-0972/. Conflict of Interest: None declared.